OpenStudy (lgbasallote):

\[y^2 = x^2\] \[\implies 2ydy - 2xdx\] \[\implies ydy = xdx\] is this implicit differentiation?

OpenStudy (anonymous):

dont start lg

OpenStudy (lgbasallote):

rainbow dps

OpenStudy (lgbasallote):

i just want to know if i have mistaught hundreds of kids

OpenStudy (anonymous):

why cant we all just get along

OpenStudy (lgbasallote):

how many of these IHMs are there

OpenStudy (anonymous):

i love math

Parth (parthkohli):

Two

OpenStudy (unklerhaukus):

i thought implicit differentiation went more like this \[y^2=x^2\] \[2y\frac{\text dy}{\text dx}=2x\]

OpenStudy (anonymous):

two hundred

OpenStudy (lgbasallote):

isn't dy same with dy/dx?

OpenStudy (unklerhaukus):

huh?

OpenStudy (lgbasallote):

dy and dy/dx are different? o.O

OpenStudy (lgbasallote):

so i really did misteach a lot of people

OpenStudy (lgbasallote):

i don't deserve to teach on OS anymore

OpenStudy (unklerhaukus):

hey we get the same result , my firs step is just different to yours ,

OpenStudy (unklerhaukus):

in you method what have you differentiated with respect to ?

Parth (parthkohli):

I don't know what the question is... but some trivia :P\[{dy \over dx}=f'(x) \]same as\[dy = f'(x)dx \]

OpenStudy (unklerhaukus):

wow you can multiply by infinitesimals

Parth (parthkohli):

Implicit differentiation is more about expressing \(y\) as an unknown function of \(x\)... so you could say the following:\[ {dy \over dx}=y'(x)\]If no info is provided.